a toy savings bank contains $17.30 consisting of nickels, dimes and quarters. the number of dimes exceeds twice the number of nickels by 3 and the number of quarters is 4 less than 5 times the number of nickels. how many quarters are in the bank?
Answer by mananth(16949)

nickels 5 x
dimes 10 (2x+3)
quarters 25 (5x-4)

Value= 1730

5x+20x+30+125x-100=1730
150x+30 =1830
150x=1830-30
150x=1800
/150
x=12 nickels 15 2x+3 dimes = 27
5x-4 = 56 quarters